OpenStudy (anonymous):

What value of x makes the equation true? -8.15 = -3.5 + x A. 4.65 B. -11.65 C. 11.65 D. -4.65

OpenStudy (anonymous):

Its B

OpenStudy (anonymous):

help with another?

OpenStudy (anonymous):

the answer is D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it D or B

OpenStudy (anonymous):

What is the value of -2.3 - q when q = 7.2? A. 9.5 B. -9.5 C. 5.1 D. -4.9

OpenStudy (anonymous):

d sorry

OpenStudy (anonymous):

What is the value of -2.3 - q when q = 7.2? A. 9.5 B. -9.5 C. 5.1 D. -4.9

OpenStudy (anonymous):

-9.5

OpenStudy (anonymous):

B

OpenStudy (anonymous):

lemme double check to confirm

OpenStudy (anonymous):

ok i have 8 more if you want to help

OpenStudy (anonymous):

ok sure and the answer i said was right

OpenStudy (anonymous):

Point G is graphed on a number line at 4. Point H is 9 units away from point G on the number line. What could be the coordinates of point H? Choose all answers that are correct. A. 13 B. -5 C. 5 D. -4

OpenStudy (anonymous):

on sec

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

I think its A and B 13 and -5

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Which equation demonstrates the opposite of a sum property? A. (13 + 7) + 3 = (7 + 13) + 3 B. -9 + 14 = 14 + (-9) C. (7 + (-2)) + 12 = 7 + ((-2) + 12) D. -(-3 + 11) = 3 + (-11)

OpenStudy (anonymous):

im not sure about this one but i can still try

OpenStudy (anonymous):

i think it might be A but i'm not completely sure

OpenStudy (anonymous):

lemme look it up

OpenStudy (anonymous):

ok never mind, do you have any more questions?

OpenStudy (anonymous):

can you guys please medal all my answers?